OpenStudy (beccab003):

Help understanding classifying! Medal to best answer. Choose the answers that correctly classify the following expression. Mark all correct answers; more than one answer may be correct. x^4+400 Monomial Difference of two squares Radical expression Sum of two squares Polynomial Complex conjugate

OpenStudy (anonymous):

Monomial: A statement with one "thing". This has two things, so, no. Difference of Two Squares: This is addition, not difference. Radical Expression: Radicals are roots like \(\sqrt{2}\), so no. Sum of Squares: \(x^4 = (x^2)^2\) and \(400 = 20^2\), so yes. Polynomial: A sum of constant multiplied powers of a variable. \(x^4 + 400 = x^4 + 400x^0\), so yes. Complex Conjugate: No.

OpenStudy (beccab003):

Thank you! :) So, Sum of Squares and Polynomial?

OpenStudy (anonymous):

Yep!